# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ Abstract
In order to explore the effect of contact with the substrate in a controllable manner, we designed an experimental system where we employ the standard formulation of a kinesin/tubulin active gel to which we add the precursors of a poly(ethylene glycol) (PEG)-based hydrogel that can be photopolymerized with UV light with the concourse of a photoinitiator [36]. We have verified that the power of the used UV radiation has no direct effect on the morphology or dynamics of the AN layer. Experiments are performed in a flow cell of 50 µm thickness, made with a superhydrophilic and a hydrophobic glass slide (Fig. 1a). The cell is filled with the active mixture and fluorinated oil to allow the formation of the active nematic at the water-oil interface. The interfaced material displays, at this stage, the well-established, seemingly chaotic flow patterns characteristic of an active nematic layer [11, 37]. We then subject the cell to a sustained (120 s) low-power-illumination. The hydrogel polymerizes, leaving a wetting film that is still capable of accommodating the microtubule-based fluid. In our experiments, we changed activity by controlling the concentration of ATP, which we kept above $300\,\mu M$ to prevent the arrest of the active nematic upon substrate polymerization. See Supplementary Material for further details on the experimental protocols.
Figure 2: Continuum simulations reproduce the two-stage transition from active nematic fluid to active biphasic fluid. (a) Schematic depiction of the simulation setup. In the beginning, simulation is carried out with only dipolar activity. At $t_{0}$, in addition to the dipolar activity, friction and quadrupolar activity are turned on to mimic the effect of turning on light in the experiments. (b)–(c) Director correlation length (rescaled by the value before $t_{0}$), $L_{C_{Q}(r)}$, and the order parameter, $S$, respectively, as a function of time (in simulation units, s.u.). The gray dashed line denotes the onset of friction and quadrupolar activity ($t_{0}^{f}=t_{0}^{q}$). Plots are averaged over 5 realizations. (d) Line integral convolution (LIC) render of snapshots of the director at different stages of transition (see Movie S2). Black regions correspond to the isotropic phase. The color of each frame corresponds to the corresponding stage indicated in panel (b). The parameters used are listed in the Supplementary Material.
@@
To investigate the physical mechanism driving the experimentally observed transition, we used a continuum model of lyotropic active nematics [38]. Here, a non-conserved scalar parameter, $\phi\in[0,1]$, is introduced to distinguish the active nematic phase driven by dipolar active stresses [33, 39, 40], from a passive, isotropic phase. When the UV light is on, the gel underneath the substrate solidifies. This leads to a friction-induced screening effect that affects the dominant forms of active stresses exerted by active entities [41]. Therefore, higher-order contributions to active stresses become important [41, 42, 43]. To account for this, we introduce quadrupolar activity to the momentum equation, together with the isotropic friction (see Supplementary Material for full model description). To replicate the experimental condition as closely as possible, these terms are switched on instantaneously at a given time after the system has reached the active turbulence state (Fig 2a). This corresponds to switching the UV light on in the experiment.
@@
Figure 3: Quadrupolar activity governs the physical mechanism of the dynamic transition. (a) LIC render of a simulated stripe of nematic phase surrounded by isotropic phase (black) with bend instability evolved under different conditions: (– –) no friction and no quadrupolar activity; (+ –) friction only; (– +) quadrupolar activity only; (+ +) both friction and quadrupolar activity. Dipolar activity is on at all times. See Movies S3–S6. (b) Correlation of director before the friction and quadrupolar activity are activated (black dashed line) and after (pink line). (c) Director correlation length (rescaled by pre-light average) and the order parameter as a function of time for high and low values of dipolar activity obtained via simulations. Plots are the average over 5 realizations. Here, $\zeta_{dipole}=\zeta_{dipole}^{0}$ (high activity) and $\zeta_{dipole}=0.1\zeta_{dipole}^{0}$ (low activity). The parameters used are listed in the Supplementary Material. (d) Same as (c) but for experiments (single realization, see Movie S7). Light power density is 3 mW cm-2, and [ATP]=1500 $\mu$M for high activity and 300 $\mu$M for low activity conditions.
@@
The model qualitatively reproduces the dynamics in the director field (see Fig. 2d and Supplementary Movie S2). Ordered regions undergo folding, breaking into smaller perpendicular domains. Afterward, the domains elongate until the active biphasic regime with smaller length scales is established. Remarkably, not only do the qualitative features mirror experimental observations, but the transition follows a two-stage dynamics as both the two-point director correlation and the order collapse (Fig. 2b–c), and the number of isotropic domains subsequently increases in agreement with experiments (Supplementary Fig. S2b). Additionally, in line with the experimental measurements, the crossover between the two stages is marked by the temporal, step-like (see Supplementary Fig. S4) increase and a peak in the director correlation length before the transition to biphasic active fluid, as well as decrease in the root-mean-squared velocity (Supplementary Fig. S5). Together, the experimental observations and continuum modeling reveal a novel two-stage process for the dynamic transition between active nematic and active biphasic fluids that is mediated by the interaction with the substrate.
@@ II Acknowledgements
We thank Kristian Thijssen for the helpful discussions. A.A. acknowledges support from the EU’s Horizon Europe research and innovation program under the Marie Sklodowska-Curie grant agreement No. 101063870 (TopCellComm). A.D. acknowledges funding from the Novo Nordisk Foundation (grant no. NNF18SA0035142 and NERD grant no. NNF21OC0068687), Villum Fonden Grant no. 29476, and the European Union via the ERC-Starting Grant PhysCoMeT, grant no. 101041418. We thank M. Pons, A. LeRoux, and G. Iruela (Universitat de Barcelona) for their assistance in the expression of motor proteins. I.V.-C. acknowledges funding from Generalitat de Catalunya through a FI-2020 PhD. Fellowship. I.V.-C., J.I.-M., and F.S. acknowledge funding from MICIU/AEI/10.13039/501100011033 (Grant No. PID2022-137713NB-C21). I.V.-C. and J.I.-M. acknowledge funding from MICIU/AEI/10.13039/501100011033 (Grant No. PDC2022-133625-I00).

### v3

@@ Abstract
In order to explore the effect of contact with the substrate in a controllable manner, we designed an experimental system where we employ the standard formulation of a kinesin/tubulin active gel to which we add the precursors of a poly(ethylene glycol) (PEG)-based hydrogel that can be photopolymerized with UV light with the concourse of a photoinitiator [36]. We have verified that the power of the used UV radiation has no direct effect on the morphology or dynamics of the AN layer. Experiments are performed in a flow cell of 50 µm thickness, made with a superhydrophilic and a hydrophobic glass slide (Fig. 1a). The cell is filled with the active mixture and fluorinated oil to allow the formation of the active nematic at the water-oil interface. The interfaced material displays, at this stage, the well-established, seemingly chaotic flow patterns characteristic of an active nematic layer [11, 37]. We then subject the cell to a sustained (120 s) low-power-illumination. The hydrogel polymerizes, leaving a wetting film that is still capable of accommodating the microtubule-based fluid. In our experiments, we changed activity by controlling the concentration of ATP, which we kept above $300\,\mu M$ to prevent the arrest of the active nematic upon substrate polymerization. See [38] for further details on the experimental protocols.
Figure 2: Continuum simulations reproduce the two-stage transition from active nematic fluid to active biphasic fluid. (a) Schematic depiction of the simulation setup. In the beginning, simulation is carried out with only dipolar activity. At $t_{0}$, in addition to the dipolar activity, friction and quadrupolar activity are turned on to mimic the effect of turning on light in the experiments. (b)–(c) Director correlation length (rescaled by the value before $t_{0}$), $L_{C_{Q}(r)}$, and the order parameter, $S$, respectively, as a function of time (in simulation units, s.u.). The gray dashed line denotes the onset of friction and quadrupolar activity ($t_{0}^{f}=t_{0}^{q}$). Plots are averaged over 5 realizations. (d) Line integral convolution (LIC) render of snapshots of the director at different stages of transition (see Movie S2). Black regions correspond to the isotropic phase. The color of each frame corresponds to the corresponding stage indicated in panel (b). The parameters used are listed in [38].
@@
To investigate the physical mechanism driving the experimentally observed transition, we used a continuum model of lyotropic active nematics [39]. Here, a non-conserved scalar parameter, $\phi\in[0,1]$, is introduced to distinguish the active nematic phase driven by dipolar active stresses [33, 40, 41], from a passive, isotropic phase. When the UV light is on, the gel underneath the substrate solidifies. This leads to a friction-induced screening effect that affects the dominant forms of active stresses exerted by active entities [42]. Therefore, higher-order contributions to active stresses become important [42, 43, 44]. To account for this, we introduce quadrupolar activity to the momentum equation, together with the isotropic friction [38]. To replicate the experimental condition as closely as possible, these terms are switched on instantaneously at a given time after the system has reached the active turbulence state (Fig 2a). This corresponds to switching the UV light on in the experiment.
@@
Figure 3: Quadrupolar activity governs the physical mechanism of the dynamic transition. (a) LIC render of a simulated stripe of nematic phase surrounded by isotropic phase (black) with bend instability evolved under different conditions: (– –) no friction and no quadrupolar activity; (+ –) friction only; (– +) quadrupolar activity only; (+ +) both friction and quadrupolar activity. Dipolar activity is on at all times. See Movies S3–S6. (b) Correlation of director before the friction and quadrupolar activity are activated (black dashed line) and after (pink line). (c) Director correlation length (rescaled by pre-light average) and the order parameter as a function of time for high and low values of dipolar activity obtained via simulations. Plots are the average over 5 realizations. Here, $\zeta_{dipole}=\zeta_{dipole}^{0}$ (high activity) and $\zeta_{dipole}=0.1\zeta_{dipole}^{0}$ (low activity). The parameters used are listed in [38]. (d) Same as (c) but for experiments (single realization, see Movie S7). Light power density is 3 mW cm-2, and [ATP]=1500 $\mu$M for high activity and 300 $\mu$M for low activity conditions.
@@
The model qualitatively reproduces the dynamics in the director field (see Fig. 2d and Supplementary Movie S2). Ordered regions undergo folding, breaking into smaller perpendicular domains. Afterwards, the domains elongate until the active biphasic regime with smaller length scales is established. Remarkably, not only do the qualitative features mirror experimental observations, but the transition follows a two-stage dynamics as both the two-point director correlation and the order collapse (Fig. 2b–c), and the number of isotropic domains subsequently increases in agreement with experiments (Supplementary Fig. S2b). Additionally, in line with the experimental measurements, the crossover between the two stages is marked by the temporal, step-like (see Supplementary Fig. S4) increase and a peak in the director correlation length before the transition to biphasic active fluid, as well as decrease in the root-mean-squared velocity (Supplementary Fig. S5). Together, the experimental observations and continuum modeling reveal a novel two-stage process for the dynamic transition between active nematic and active biphasic fluids that is mediated by the interaction with the substrate.
@@ II Acknowledgements
We thank Kristian Thijssen for the helpful discussions. A.A. acknowledges support from the EU’s Horizon Europe research and innovation program under the Marie Sklodowska-Curie grant agreement No. 101063870 (TopCellComm). A.D. acknowledges funding from the Novo Nordisk Foundation (grant no. NNF18SA0035142 and NERD grant no. NNF21OC0068687), Villum Fonden Grant no. 29476, and the European Union via the ERC-Starting Grant PhysCoMeT, grant no. 101041418. We thank M. Pons, A. LeRoux, and G. Iruela (Universitat de Barcelona) for their assistance in the expression of motor proteins. I.V.-C. acknowledges funding from Generalitat de Catalunya through a FI-2020 PhD. Fellowship. I.V.-C., J.I.-M., and F.S. acknowledge funding from MICIU/AEI/10.13039/501100011033 (Grant No. PID2022-137713NB-C21). I.V.-C. and J.I.-M. acknowledge funding from MICIU/AEI/10.13039/501100011033 (Grant No. PDC2022-133625-I00). The authors are indebted to the Brandeis University MRSEC Biosynthesis facility, supported by NSF DMR-MRSEC 2011846, for providing the tubulin.